# Quintessential early dark energy

### v2

@@ Abstract
In this paper, we introduce a unified model of early and late dark energy. We call it quintessential early dark energy model in which early and late dark energy are explained by a single scalar field. In other words two different energy scales are related by a single scalar field potential. To achieve this we introduce the modified steep exponential potential. This potential has a hilltop nature during the early time which consists of a flat region followed by a steep region. This nature of the potential plays a crucial role in achieving early dark energy solution. During recent time, the potential can almost mimic the cosmological constant which can result into late time acceleration. But, at the perturbation level the potential shows significant difference with the $\Lambda$CDM model. We also constrain and compare the models for steep exponential, modified steep exponential, axionlike and power law potentials by using the available background cosmological data from CMB, BAO (including DESI DR1 2024), supernovae (Pantheon$+$, DESY5 and Union3) and Hubble parameter measurements. Even after the presence of required EDE solution in all four potentials we don’t get any significant improvement in the value of $H_{0}$. The maximum improvement we get in the present value of Hubble parameter compared to the standard $\Lambda$CDM model is for the axionlike potential. For other potentials the constraints are similar to the $\Lambda$CDM model. We also see that the data prefers $\Lambda$CDM model over the considered scalar field models at least for the data combinations with Pantheon$+$ and Union3.
@@ 1 Introduction
As we have discussed about the importance of the steep nature of the potential for the EDE solution it also makes the scalar field energy density decay very fast or scale the background [37, 58]. Both the cases do not give rise to late time acceleration. Generally, if we have EDE solution then the late time dark energy (LDE) should be achieved either by adding a CC or an another potential. So, it would be interesting if we can unify the EDE and LDE with a single scalar field potential. We explore this possibility in this paper and we modify the SE potential such a way so that it gives EDE at higher redshift and LDE during the recent time. The modified steep exponential (MSE) potential behaves like the SE potential during early times and from recent past it behaves almost like a CC and thereby results LDE. While producing LDE the scalar field rolls slowly. This slowly rolling scalar field is known as quintessence [57, 59, 60]. Following the concept of quintessential inflation [61, 62, 63, 64] in which inflation and LDE are described through a single scalar field we call the scenario under consideration where a single scalar field results both EDE and LDE as Quintessential EDE. It is important to note here that in the scenarios like quintessential inflation or quintessential EDE there should be one energy scale associated with the scalar field potential which can be related to another energy scale and thereby resulting the unification of the two energy scales. If any potential has two energy scales [65, 66, 67] then it is trivial to represent the cosmic history of two different scales with that potential but it is not unification of two energy scales. In this regard, it is also important to mention that the potentials which can give rise to scaling solutions may not be able to give sufficient amount of EDE followed by a small amount of scalar field energy density during matter dominated era and therefor may not alleviate the $H_{0}$ tension for a canonical scalar field [65, 68].
@@ 2 Background equations
$\mathcal{S}=\int\mathrm{d}^{4}x\sqrt{-\mathrm{g}}\Bigl{[}\frac{M_{\textrm{Pl}}^{2}}{2}R-\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi-V(\phi)\Bigr{]}+\mathcal{S}_{\mathrm{m}}+\mathcal{S}_{\mathrm{r}}\,,$ (2.1)
@@
Varying the action (2.1) with respect to (w.r.t.) the metric $g_{\mu\nu}$ gives the Einstein’s field equation
$M_{\textrm{Pl}}^{2}G_{\mu\nu}=T_{(\mathrm{m})\mu\nu}+T_{(\mathrm{r})\mu\nu}+T_{(\phi)\mu\nu}\,,$ (2.2)
@@
$T_{(\phi)\mu\nu}=$ $\phi_{;\mu}\phi_{;\nu}-\frac{1}{2}\mathrm{g}_{\mu\nu}(\nabla\phi)^{2}-\mathrm{g}_{\mu\nu}V(\phi)\,.$ (2.3)
The equation of motion of the scalar field can be calculated by varying the action (2.1) w.r.t. the scalar field $\phi$ an it is given by
$\Box\phi-V_{\phi}(\phi)=0\,,$ (2.4)
@@
$3H^{2}M_{\textrm{Pl}}^{2}$ $=$ $\rho_{\mathrm{m}}+\rho_{\mathrm{r}}+\rho_{\phi}\,$ (2.6)
$\left(2\dot{H}+3H^{2}\right)M_{\textrm{Pl}}^{2}$ $=$ $-\frac{1}{3}\rho_{\mathrm{r}}-p_{\phi}\,,$ (2.7)
@@
$\ddot{\phi}+3H\dot{\phi}+\frac{\mathrm{d}V}{\mathrm{d}\phi}=0\,.$ (2.10)
@@
$w_{\text{eff}}$ $=$ $-\left(1+\frac{2}{3}\frac{\dot{H}}{H^{2}}\right)\;,$ (2.11)
$w_{\phi}$ $=$ $\frac{\frac{1}{2}\dot{\phi}^{2}-V(\phi)}{\frac{1}{2}\dot{\phi}^{2}+V(\phi)}\;.$ (2.12)
@@
$\lambda$ $=$ $-M_{\textrm{Pl}}\frac{V^{\prime}(\phi)}{V}\;,$ (2.13)
$\Gamma$ $=$ $\frac{V^{\prime\prime}(\phi)V(\phi)}{V^{\prime}(\phi)^{2}}\;.$ (2.14)
@@ 3 Steep Exponential Potential
where $V_{0}$, $\lambda$ and $n$ are constants. Potential (3.1) is a steep exponential (SE) potential. Introduced in [58], SE is well studied [70, 71, 72, 73, 74, 75, 76, 77, 78] especially in the field of quintessential inflation models [62] where inflation and late time acceleration are achieved by a single scalar field. Since the potential (3.1) is very steep we can’t get late time acceleration only with the potential which eventually leads to a scaling solution [58]. To achieve late time acceleration we need to consider an interaction of the scalar field with the massive neutrino which plays a crucial roll to achieve late time acceleration [58].
@@
Figure 1: Both the figures the solid blue, dashed green and dotted red lines represent energy densities, normalised by present critical density, of scalar field, matter and radiation respectively. Left: Initial conditions are $\phi_{i}=1.4$, $\phi_{i}^{\prime}=0.15$ with parameters $n=4$, $\mu=0.606$ and $V_{0}=5.58\times 10^{12}$. Right: Initial conditions are $\phi_{i}=6$, $\phi_{i}^{\prime}=0.15$ with parameters $n=4$, $\mu=0.012$ and $V_{0}=1.8\times 10^{17}$.
For the SE potential (3.1) the functions $\lambda$ and $\Gamma$ are given by [58]
$\lambda$ $=$ $\mu n\left(\frac{\phi}{M_{\textrm{Pl}}}\right)^{n-1}\,,$ (3.2)
$\Gamma$ $=$ $1-\frac{n-1}{\mu n}\frac{M_{\textrm{Pl}}^{n}}{\phi^{n}}\,.$ (3.3)
@@
For $n=1$ the potential (3.1) reduces to the exponential potential with $\lambda=$constant and $\Gamma=1$. For $n>1$ and $\phi>0$ we see, form the Eqs. (3.2) and (3.3), that $\lambda$ is ever increasing but $\Gamma$ approaches $1$ as $\phi$ increases which makes the potential effectively an exponential potential with very large slope. These natures of $\lambda$ and $\Gamma$ for the SE potential gives an interesting dynamics of the scalar field. For $n>1$ and around $\phi=0$ the potential is very flat and we can have slow roll in that region. This flat region is followed by an steep region for $\phi>0$ where the scalar field will initially roll fast giving rise to $w_{\phi}=1$ and $\rho_{\phi}\sim a^{-6}$. Because of this sharp fall in the energy density the scalar field will experience a large Hubble damping due to the background and the scalar field will frozen giving rise to cosmological constant like behaviour in the scalar field dynamics. During this frozen period $\rho_{\phi}$ will increase and when it becomes comparable to the background energy density the scalar field again starts rolling. Now, after exiting from the frozen period the dynamics of the scalar field will be determined by the value of the field $\phi$. If $\phi$ is sufficiently large then $\Gamma$ will be close to $1$ which will lead to a scaling behaviour and if $\phi$ is not large enough then due to the steep nature of the potential the scalar field will repeat the previous dynamics [58] which has been depicted in the Fig. 1. The dynamics shown in the left figure of Fig. 1 can be useful to achieve early dark energy. For early dark energy the scalar field should be frozen before the matter radiation equality and start evolving from around matter radiation equality. This behaviour will inject a small fraction of dark energy during the matter radiation equality which can alleviate the Hubble tension to some extent [11, 38]. We shall discuss this in Sec. 5.
@@
$\Gamma=1$ and $\lambda>\sqrt{3}$ gives rise to scaling behaviour [33] in the scalar field dynamics where the scalar field energy density follows the background energy density. Scaling behaviour can not give rise to late time acceleration. So if a potential has large slope $\lambda$ and $\Gamma=1$ then it will give rise to scaling behaviour but late time acceleration can not be achieved. Now, to get the both scaling behaviour and the late time acceleration we have to exit from the scaling behaviour in the recent past i.e., $\Gamma$ should go away from 1 and $\lambda$ should approach zero. One way to achieve this is to add a cosmological constant in the SE potential (3.1). But here we do not want to add a cosmological constant explicitly and rather we modify the SE potential such a way so that it gives rise to a cosmological constant like behaviour from the recent past. We have discussed this in the next section.
@@ 4 Modified Steep Exponential Potential
Figure 2: Nature of the MSE potential (4.1) for different values of $n$ and $\gamma$.
As we have discussed in the previous section that the SE potential (3.1) can give rise to scaling behaviour but not late time acceleration. The steep nature of the SE potential can be useful to achieve EDE solutions. Now,to get late time acceleration from a steep exponential potential while retaining the properties of scaling-like behaviour and possible EDE solutions we consider the following modified steep exponential potential (MSE)
$V(\phi)$ $=$ $V_{0}\mathrm{e}^{-\gamma\frac{(\phi/M_{\textrm{Pl}})^{n}}{\phi_{0}+(\phi/Mpl)^{n}}}\,,$ (4.1)
where $V_{0}$, $\phi_{0}$, $\gamma$ and $n$ are constants. For both the potentials (3.1) and (4.1), $V_{0}$ fixes the energy scale of the potential. Increasing the values of $n$ increases the flatness of the potentials around $\phi=0$. It also increases the steepness of the potential. $\gamma$ is related to the slope of the potential and changes the steepness. We should note that for $\phi/M_{\textrm{Pl}}\ll\phi_{0}$ the MSE potential (4.1) reduces to the potential (3.1) with $\mu=\gamma/\phi_{0}$. But in the potential (4.1) we do not need to add any CC or an interaction as for $\phi/M_{\textrm{Pl}}\gg\phi_{0}$ the potential will be very flat and can give late time acceleration. Dependence of the potential (4.1) on the $n$ and $\gamma$ has been shown in Fig. 2.
@@ 4.1 Relating different energy scales by MSE: Unifying EDE and LDE
The MSE potential (4.1) can be useful to relate two different energy scales with a single canonical scalar field without considering any interaction as we have mentioned above. For small positive field values i.e., $\phi/M_{\textrm{Pl}}\ll\phi_{0}$ the potential (4.1) becomes
$V(\phi)$ $=$ $V_{0}\mathrm{e}^{-\mu(\phi/M_{\textrm{Pl}})^{n}}\,,$ (4.2)
$\mu$ $=$ $\gamma/\phi_{0}\,,$ (4.3)
which has the same form as the SE potential (3.1). For large field values let’s write the MSE potential (4.1) in the following form
$V(\phi)$ $=$ $V_{0}\mathrm{e}^{-\gamma(1+\phi_{0}/(\phi/M_{\textrm{Pl}})^{n})^{-1}}\,.$ (4.4)
@@
$V_{0}\mathrm{e}^{-\gamma}=V_{1}\,,$ (4.5)
i.e, the MSE potential (4.1) becomes a constant potential and one can calculate $V_{1}$ once we choose $V_{0}$ and $\gamma$ and thereby we relate two energy scales $V_{0}$ and $V_{1}$ by a single scalar field. So, the potential (4.1) can give late time acceleration without considering an interaction between scalar field and massive neutrinos [58] i.e., we can achieve the unified scenarios only from a canonical scalar field with the potential (4.1). The last equation gives us
$\gamma=-\ln\frac{V_{1}}{V_{0}}\,.$ (4.6)
@@
$\phi_{1}/M_{\textrm{Pl}}$ $=$ $\phi_{0}^{1/n}\left(\frac{\gamma}{\ln(V_{0}/V_{1})}-1\right)^{-1/n}\,.$ (4.7)
While the Eq.(4.6) gives an estimate about the value of $\gamma$ needed to relate $V_{1}$ with $V_{0}$ the Eq. (4.7) gives the bound on the value of $\gamma$ which is given by
$\gamma>\ln(V_{0}/V_{1})\,.$ (4.8)
Since for large field values the potential becomes very flat it can behave like a CC and by choosing $V_{1}=\rho_{\rm DE}$, where $\rho_{\rm DE}$ is the dark energy density, we can relate the dark energy scale with some higher energy scale. E.g., if we are interested to relate dark energy scale with that of the inflation scale then $V_{0}^{1/4}\sim 10^{16}$eV which gives us $\gamma\approx 255$. Similarly we can choose $V_{0}^{1/4}\sim 10^{-10}$eV which is around matter-radiation equality we get $\gamma\approx 18$ which basically unifies the early and late dark energy. Important thing to notice is that the values of the parameter $\gamma$ are natural and we do not have to deal with the unnatural fine tuning of the dark energy scale.

### v3

@@ Abstract
We introduce a unified model of early and late dark energy. We call it quintessential early dark energy model where early and late dark energy are explained by a single scalar field i.e., two different energy scales are related by a single scalar field potential. To achieve this we introduce the modified steep exponential potential, which is chosen phenomenologically. This potential has a hilltop nature during the early time which consists of a flat region followed by a steep region. This nature of the potential plays a crucial role in achieving early dark energy solution. During recent time, the potential can almost mimic the cosmological constant which can result into late time acceleration. But, at the perturbation level the potential shows significant difference with the $\Lambda$CDM model. We also constrain and compare the models for steep exponential, modified steep exponential, axionlike and power law potentials by using the available background cosmological data from CMB, BAO (including DESI DR1 2024), supernovae (Pantheon$+$, DESY5 and Union3) and Hubble parameter measurements. Even after the presence of required EDE solution in all four potentials we don’t get any significant improvement in the value of $H_{0}$. The maximum improvement we get in the present value of Hubble parameter compared to the standard $\Lambda$CDM model is for the axionlike potential. For other potentials the constraints are similar to the $\Lambda$CDM model. We also see that the data prefers $\Lambda$CDM model over the considered scalar field models at least for the data combinations with Pantheon$+$ and Union3.
keywords:
Unified dark energy models, Early and late dark energy, Cosmological tensions, Cosmological observations
journal: Physics of the Dark Universe
[…]
[1]organization=Department of Physics, Jamia Millia Islamia,city=New Delhi, postcode=110025, country=India
[…]
[2]organization=Department of Astronomy, Astrophysics & Space Engineering, Indian Institute of Technology,city=Indore, postcode=453552, country=India
@@ 1 Introduction
As we have discussed about the importance of the steep nature of the potential for the EDE solution it also makes the scalar field energy density decay very fast or scale the background [37, 58]. Both the cases do not give rise to late time acceleration. Generally, if we have EDE solution then the late time dark energy (LDE) should be achieved either by adding a CC or an another potential. So, it would be interesting if we can unify the EDE and LDE with a single scalar field potential. We explore this possibility in this paper and we modify the SE potential such a way so that it gives EDE at higher redshift and LDE during the recent time. The modified steep exponential (MSE) potential, chosen phenomenologically, behaves like the SE potential during early times and from recent past it behaves almost like a CC and thereby results LDE. While producing LDE the scalar field rolls slowly. This slowly rolling scalar field is known as quintessence [57, 59, 60]. Following the concept of quintessential inflation [61, 62, 63, 64] in which inflation and LDE are described through a single scalar field we call the scenario under consideration where a single scalar field results both EDE and LDE as Quintessential EDE. It is important to note here that in the scenarios like quintessential inflation or quintessential EDE there should be one energy scale associated with the scalar field potential which can be related to another energy scale and thereby resulting the unification of the two energy scales. If any potential has two energy scales [65, 66, 67] then it is trivial to represent the cosmic history of two different scales with that potential but it is not unification of two energy scales. In this regard, it is also important to mention that the potentials which can give rise to scaling solutions may not be able to give sufficient amount of EDE followed by a small amount of scalar field energy density during matter dominated era and therefor may not alleviate the $H_{0}$ tension for a canonical scalar field [65, 68].
@@ 2 Background equations
$\mathcal{S}=\int\mathrm{d}^{4}x\sqrt{-\mathrm{g}}\Bigl{[}\frac{M_{\textrm{Pl}}^{2}}{2}R-\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi-V(\phi)\Bigr{]}+\mathcal{S}_{\mathrm{m}}+\mathcal{S}_{\mathrm{r}}\,,$ (1)
@@
Varying the action (1) with respect to (w.r.t.) the metric $g_{\mu\nu}$ gives the Einstein’s field equation
$M_{\textrm{Pl}}^{2}G_{\mu\nu}=T_{(\mathrm{m})\mu\nu}+T_{(\mathrm{r})\mu\nu}+T_{(\phi)\mu\nu}\,,$ (2)
@@
$T_{(\phi)\mu\nu}=$ $\phi_{;\mu}\phi_{;\nu}-\frac{1}{2}\mathrm{g}_{\mu\nu}(\nabla\phi)^{2}-\mathrm{g}_{\mu\nu}V(\phi)\,.$ (3)
The equation of motion of the scalar field can be calculated by varying the action (1) w.r.t. the scalar field $\phi$ an it is given by
$\Box\phi-V_{\phi}(\phi)=0\,,$ (4)
@@
$3H^{2}M_{\textrm{Pl}}^{2}$ $=$ $\rho_{\mathrm{m}}+\rho_{\mathrm{r}}+\rho_{\phi}\,$ (6)
$\left(2\dot{H}+3H^{2}\right)M_{\textrm{Pl}}^{2}$ $=$ $-\frac{1}{3}\rho_{\mathrm{r}}-p_{\phi}\,,$ (7)
@@
$\ddot{\phi}+3H\dot{\phi}+\frac{\mathrm{d}V}{\mathrm{d}\phi}=0\,.$ (10)
@@
$w_{\text{eff}}$ $=$ $-\left(1+\frac{2}{3}\frac{\dot{H}}{H^{2}}\right)\;,$ (11)
$w_{\phi}$ $=$ $\frac{\frac{1}{2}\dot{\phi}^{2}-V(\phi)}{\frac{1}{2}\dot{\phi}^{2}+V(\phi)}\;.$ (12)
@@
$\lambda$ $=$ $-M_{\textrm{Pl}}\frac{V^{\prime}(\phi)}{V}\;,$ (13)
$\Gamma$ $=$ $\frac{V^{\prime\prime}(\phi)V(\phi)}{V^{\prime}(\phi)^{2}}\;.$ (14)
@@ 3 Steep Exponential Potential
where $V_{0}$, $\lambda$ and $n$ are constants. Potential (15) is a steep exponential (SE) potential. Introduced in [58], SE is well studied [70, 71, 72, 73, 74, 75, 76, 77, 78] especially in the field of quintessential inflation models [62] where inflation and late time acceleration are achieved by a single scalar field. Since the potential (15) is very steep we can’t get late time acceleration only with the potential which eventually leads to a scaling solution [58]. To achieve late time acceleration we need to consider an interaction of the scalar field with the massive neutrino which plays a crucial roll to achieve late time acceleration [58].
@@
Figure 1: In both the figures the solid blue, dashed green and dotted red lines represent energy densities, normalised by present critical density, of scalar field, matter and radiation respectively. Left: Initial conditions are $\phi_{i}=1.4$, $\phi_{i}^{\prime}=0.15$ with parameters $n=4$, $\mu=0.606$ and $V_{0}=5.58\times 10^{12}$. Right: Initial conditions are $\phi_{i}=6$, $\phi_{i}^{\prime}=0.15$ with parameters $n=4$, $\mu=0.012$ and $V_{0}=1.8\times 10^{17}$.
For the SE potential (15) the functions $\lambda$ and $\Gamma$ are given by [58]
$\lambda$ $=$ $\mu n\left(\frac{\phi}{M_{\textrm{Pl}}}\right)^{n-1}\,,$ (16)
$\Gamma$ $=$ $1-\frac{n-1}{\mu n}\frac{M_{\textrm{Pl}}^{n}}{\phi^{n}}\,.$ (17)
@@
For $n=1$ the potential (15) reduces to the exponential potential with $\lambda=$constant and $\Gamma=1$. For $n>1$ and $\phi>0$ we see, form the Eqs. (16) and (17), that $\lambda$ is ever increasing but $\Gamma$ approaches $1$ as $\phi$ increases which makes the potential effectively an exponential potential with very large slope. These natures of $\lambda$ and $\Gamma$ for the SE potential gives an interesting dynamics of the scalar field. For $n>1$ and around $\phi=0$ the potential is very flat and we can have slow roll in that region. This flat region is followed by an steep region for $\phi>0$ where the scalar field will initially roll fast giving rise to $w_{\phi}=1$ and $\rho_{\phi}\sim a^{-6}$. Because of this sharp fall in the energy density the scalar field will experience a large Hubble damping due to the background and the scalar field will frozen giving rise to cosmological constant like behaviour in the scalar field dynamics. During this frozen period $\rho_{\phi}$ will increase and when it becomes comparable to the background energy density the scalar field again starts rolling. Now, after exiting from the frozen period the dynamics of the scalar field will be determined by the value of the field $\phi$. If $\phi$ is sufficiently large then $\Gamma$ will be close to $1$ which will lead to a scaling behaviour and if $\phi$ is not large enough then due to the steep nature of the potential the scalar field will repeat the previous dynamics [58] which has been depicted in the Fig. 1. The dynamics shown in the left figure of Fig. 1 can be useful to achieve early dark energy. For early dark energy the scalar field should be frozen before the matter radiation equality and start evolving from around matter radiation equality. This behaviour will inject a small fraction of dark energy during the matter radiation equality which can alleviate the Hubble tension to some extent [11, 38]. We shall discuss this in Sec. 5.
@@
$\Gamma=1$ and $\lambda>\sqrt{3}$ gives rise to scaling behaviour [33] in the scalar field dynamics where the scalar field energy density follows the background energy density. Scaling behaviour can not give rise to late time acceleration. So if a potential has large slope $\lambda$ and $\Gamma=1$ then it will give rise to scaling behaviour but late time acceleration can not be achieved. Now, to get the both scaling behaviour and the late time acceleration we have to exit from the scaling behaviour in the recent past i.e., $\Gamma$ should go away from 1 and $\lambda$ should approach zero. One way to achieve this is to add a cosmological constant in the SE potential (15). But here we do not want to add a cosmological constant explicitly and rather we modify the SE potential such a way so that it gives rise to a cosmological constant like behaviour from the recent past. We have discussed this in the next section.
@@ 4 Modified Steep Exponential Potential
Figure 2: Nature of the MSE potential (18) for different values of $n$ and $\gamma$.
As we have discussed in the previous section that the SE potential (15) can give rise to scaling behaviour but not late time acceleration. The steep nature of the SE potential can be useful to achieve EDE solutions. Now,to get late time acceleration from a steep exponential potential while retaining the properties of scaling-like behaviour and possible EDE solutions we consider the following modified steep exponential potential (MSE)
$V(\phi)$ $=$ $V_{0}\mathrm{e}^{-\gamma\frac{(\phi/M_{\textrm{Pl}})^{n}}{\phi_{0}+(\phi/Mpl)^{n}}}\,,$ (18)
where $V_{0}$, $\phi_{0}$, $\gamma$ and $n$ are constants. For both the potentials (15) and (18), $V_{0}$ fixes the energy scale of the potential. Increasing the values of $n$ increases the flatness of the potentials around $\phi=0$. It also increases the steepness of the potential. $\gamma$ is related to the slope of the potential and changes the steepness. We should note that for $\phi/M_{\textrm{Pl}}\ll\phi_{0}$ the MSE potential (18) reduces to the potential (15) with $\mu=\gamma/\phi_{0}$. But in the potential (18) we do not need to add any CC or an interaction as for $\phi/M_{\textrm{Pl}}\gg\phi_{0}$ the potential will be very flat and can give late time acceleration. Dependence of the potential (18) on the $n$ and $\gamma$ has been shown in Fig. 2.
@@ 4.1 Relating different energy scales by MSE: Unifying EDE and LDE
The MSE potential (18) can be useful to relate two different energy scales with a single canonical scalar field without considering any interaction as we have mentioned above. For small positive field values i.e., $\phi/M_{\textrm{Pl}}\ll\phi_{0}$ the potential (18) becomes
$V(\phi)$ $=$ $V_{0}\mathrm{e}^{-\mu(\phi/M_{\textrm{Pl}})^{n}}\,,$ (19)
$\mu$ $=$ $\gamma/\phi_{0}\,,$ (20)
which has the same form as the SE potential (15). For large field values let’s write the MSE potential (18) in the following form
$V(\phi)$ $=$ $V_{0}\mathrm{e}^{-\gamma(1+\phi_{0}/(\phi/M_{\textrm{Pl}})^{n})^{-1}}\,.$ (21)
@@
$V_{0}\mathrm{e}^{-\gamma}=V_{1}\,,$ (22)
i.e, the MSE potential (18) becomes a constant potential and one can calculate $V_{1}$ once we choose $V_{0}$ and $\gamma$ and thereby we relate two energy scales $V_{0}$ and $V_{1}$ by a single scalar field. So, the potential (18) can give late time acceleration without considering an interaction between scalar field and massive neutrinos [58] i.e., we can achieve the unified scenarios only from a canonical scalar field with the potential (18). The last equation gives us
$\gamma=-\ln\frac{V_{1}}{V_{0}}\,.$ (23)
@@
$\phi_{1}/M_{\textrm{Pl}}$ $=$ $\phi_{0}^{1/n}\left(\frac{\gamma}{\ln(V_{0}/V_{1})}-1\right)^{-1/n}\,.$ (24)
While the Eq.(23) gives an estimate about the value of $\gamma$ needed to relate $V_{1}$ with $V_{0}$ the Eq. (24) gives the bound on the value of $\gamma$ which is given by
$\gamma>\ln(V_{0}/V_{1})\,.$ (25)
Since for large field values the potential becomes very flat it can behave like a CC and by choosing $V_{1}=\rho_{\rm DE0}$, where $\rho_{\rm DE0}$ is the dark energy density at present, we can relate the dark energy scale with some higher energy scale. E.g., if we are interested to relate dark energy scale with that of the inflation scale then $V_{0}^{1/4}\sim 10^{16}$eV which gives us $\gamma\approx 255$. Similarly we can choose $V_{0}^{1/4}\sim 10^{-10}$eV which is around matter-radiation equality we get $\gamma\approx 18$ which basically unifies the early and late dark energy. Important thing to notice is that the values of the parameter $\gamma$ are natural and we do not have to deal with the unnatural fine tuning of the dark energy scale.